OpenStudy (anonymous):

Does this f(x)=x3−12x2+28x−9 change to f(−x)=−x3+12x2−28x−9 ?

OpenStudy (anonymous):

or am i doing it wrong?

OpenStudy (anonymous):

And this one... this f(x)=x4−x3+7x2−9x−18 to this f(−x)=x4+x3+7x2+9x−18

OpenStudy (anonymous):

2nd one is absolutely right.

OpenStudy (anonymous):

thats what i put didnt i?

OpenStudy (anonymous):

odd powers doesnt affect negative sign but even power do change.

OpenStudy (anonymous):

i wrote wrong expression.sorry my fault.but I hope u understood the method.

OpenStudy (anonymous):

what? isnt it odd exponents cause the x to be replaced with -x? and even exponents stay the same?

OpenStudy (anonymous):

@heybhai

OpenStudy (jdoe0001):

hmm, if the exponent is odd, the -x will change the term, if it's even, it won't

OpenStudy (anonymous):

okay well he told me differently but he still gave me the same function i wrote and said i was wrong. im confused

OpenStudy (anonymous):

is the first one converted right?

OpenStudy (anonymous):

(-x)^3-12(-x)^2+28(-x)-9 -x^3-12x-28x-9 This is right one

OpenStudy (jdoe0001):

\(\bf f(x)=x^3−12x^2+28x-9\\ \text{change to }\\ f(-x)=(-x)^3−12(-x)^2+28(-x)-9\\\ \implies f(x)=-x^3−12x^2-28x-9\)

OpenStudy (anonymous):

but 12x^2 has an even exponent....

OpenStudy (jdoe0001):

12x^2 is negative from the beginning so it keeps whatever sign it had before

OpenStudy (jdoe0001):

if the exponent is even, the -x doesn't change the sign

OpenStudy (anonymous):

oh okay i think i just over looked that. thanks for clearing it up(:

OpenStudy (jdoe0001):

\(\bf -12(-x)^2 \implies -12(-x)(-x) \implies -12(+x)^2\)